LOFAR Newsletters June 2015

Distributed on 22 June 2015.

 

Array status:

----------------------

 

- 38 stations operational in the Netherlands: 24 core stations and 14 remote stations.

- 9 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609.

- 3 new international stations are being built in Poland and should become part of the array by the end of 2015.

- RS305 suffered a power outage on Friday 5 June. Further investigations have shown that many LBA elements were also damaged and that this was caused by lightning. Maintenance is in progress to repair them.

- RFI sources nearby DE609 affect the quality of the LBA data. Investigation is in progress to solve this issue.

- The overview of non-operational antenna elements for LBA and HBA is available here. The maintenance season is in progress.

- Station calibration: new calibration tables for mode 1 have been installed at several stations last week. Mode 3 caltables will be ready by the end of this week and will be installed as soon as the observing schedule allows.  Once the maintenance of the stations is finalized, calibration will be performed again for all modes.

 

Observing System Status

------------------------------------ 

 

- In the period 18-29 May, a successful reconfiguration of the LOFAR network infrastructure was performed. With this, the LOFAR network is prepared for the future by allowing the extension of LOFAR with new compute systems, such as DRAGNET and CEP4, as well as new LOFAR stations, such as the Polish stations.

- The observing system was mostly stable during the period Apr - Jun 2015. However, a few failures at COBALT nodes, as well as (not announced) outages at the international station lines, warm temperatures, and the instability of processing and ingest systems caused delays in the exploitation of the Cycle 3&4 observing programmes and in digesting data out of CEP2.

- Lower ingest speeds towards the Juelich LTA location are experienced during ILT runs involving all international stations - investigations are in progress at Juelich to try to fix this issue.

 

Software Development Status (H. Holties):

------------------------------------------

In May, we welcomed Jorrit Schaap as the new software developer at the ASTRON Radio Observatory. Initially, Jorrit will focus on LOFAR LTA ingest related topics. The June LOFAR Software Release 2.11, planned to be rolled out on June 29th, includes many bugfixes and stability improvements as well as the introduction of casacore 2.0.1 on all systems, a new message broker enabled version of the software handling data transfers from CEP to the LTA, and addressing a number of relief issues in MoM. Other efforts included support for characterization of the SelfCal pipeline, support for the LOFAR network reconfiguration, support for ingesting EOR data products into the LTA, and introducing message broker communication to the pipeline framework. The latter will likely be incorporated in an intermediate coming software release.

 

 

CITT Status (G. Heald):

-----------------

The Calibration and Imaging Tiger Team (CITT) is nearing the end of its 2-year project timeline and is therefore wrapping up its activities. We have just concluded a productive work week in Dwingeloo (8-12 June) that allowed us to focus the development of our direction-dependent calibration and imaging pipeline, scheduled for early community testing in the very near future. This pipeline is being built upon a flexible upgrade of the LOFAR pipeline framework itself and will allow straightforward incorporation of new software advances as they become available, even after the end of the CITT project. Early astronomical testing of the new version of the awimager (including fundamental improvements in accuracy and speed) was also carried out and produced very promising results. As we work toward concluding these development activities we are also ensuring that corresponding documentation will be available.

 

 

MSSS status (G. Heald):

------------------

The MSSS workshop held at ASTRON during 4-5 June was very productive. In advance, a preliminary version of the unified all-sky HBA catalog (including ~150,000 sources) was released to a group of researchers including the workshop participants, using our VO-capable web server. This first internal release was used for initial testing and characterisation that was reported during the meeting. As expected, some issues were identified with the data products. A revised version of the initial catalog is now in production and will be ready for testing soon. As we verify the astronomical quality of the MSSS-HBA catalog we are also preparing for a first collection of MSSS science papers to coincide with the forthcoming public release of the images and catalog. Continuing progress toward those milestones will continue to be reported in the LOFAR Status Meetings (LSM) and in this newsletter.

 

Observing Programmes and miscellaneous items:

--------------------------------------------------------------------

- Cycle 2 observing programme: 91% complete. Closed on May 15th 2015.

- Cycle 3 observing programme: 89 % complete - Remaining 11% is being observed with second priority during Cycle 4.

- Cycle 4 observing programme: 15% complete.

- 36 projects were awarded observing and processing time for Cycle 4 at the PC meeting on May 6-7 2015. The allocations are detailed here. The Cycle 4 observing schedule is available here.

- A polarization leakage affected LOFAR beamformed data recorded between July 12th 2014 and January 21th 2015. The PIs of all affected projects have been notified. A full description of the issue is available here.

- The 2nd LOFAR Users Meeting and the 2015 LOFAR Science Community workshop took place in Assen on June 1-3 2015.  The meetings brought together 90 members of the LOFAR science community.  The presentations given at these events are available on the meetings website.

 

CEP news:

----------------

 

- CEP2: disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends.

- CEP3:

- lof019 experienced a corrupted file system on May 18 2015.

- Nodes seem to be under-used or not used at all during the originally awarded period.  Monitoring of node usage is active now. These statistics will be used to evaluate requests for extensions. 

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

- Next LSM's : 24/06, 08/07, 22/07, 19/08, 02/09

(note: all presentations given at the LSM are available here).

- Next Stop days: 07/07, 04/08, 01/09.

- Next software roll outs: 29/06, 07/09.

- Cycle 5 proposal call: beginning of July 2015.

- Cycle 5 proposal submission deadline: beginning of September 2015.

- Next LOFAR bulletin: August 2015.

 

Design: Kuenst.    Development: Dripl.    © 2020 ASTRON